# Deformable Endoscopic Tissues Reconstruction with Gaussian Splatting

### v1

@@ Abstract
Surgical 3D reconstruction is a critical area of research in robotic surgery, with recent works adopting variants of dynamic radiance fields to achieve success in 3D reconstruction of deformable tissues from single-viewpoint videos. However, these methods often suffer from time-consuming optimization or inferior quality, limiting their adoption in downstream tasks. Inspired by 3D Gaussian Splatting, a recent trending 3D representation, we present EndoGS, applying Gaussian Splatting for deformable endoscopic tissue reconstruction. Specifically, our approach incorporates deformation fields to handle dynamic scenes, depth-guided supervision to optimize 3D targets with a single viewpoint, and a spatial-temporal weight mask to mitigate tool occlusion. As a result, EndoGS reconstructs and renders high-quality deformable endoscopic tissues from a single-viewpoint video, estimated depth maps, and labeled tool masks. Experiments on DaVinci robotic surgery videos demonstrate that EndoGS achieves superior rendering quality. Code is available at https://github.com/HKU-MedAI/EndoGS.
@@ 1 Introduction
Reconstruction high-quality deformable tissues from endoscopic videos is a significant but challenging task, facilitating downstream tasks like surgical AR, education, and robot learning [3, 22, 23]. Earlier attempts [13, 20, 24, 32, 33] adopt depth estimation to achieve great success in endoscopic reconstruction, but these methods still struggle to produce realistic 3D reconstruction due to two key issues. First, non-rigid deformations with sometimes large movements, which require practical dynamic scene reconstruction, hinders the adaptation of those techniques. Second, tool occlusion happens in single-viewpoint videos, producing difficulties in learning affected parts with limited information.
@@
While [9, 12] have proposed frameworks combining tool masking, stereo depth estimation, and sparse warp field [6, 16] for single-viewpoint 3D reconstruction, they are still prone to failure in the presence of dramatic non-topological deformable tissues changes. Neural Radiance Fields (NeRFs) [15] have shown great potentials in 3D representations, and methods based on variants of dynamic radiance filed [18, 5] have become representative works in deformable tissues reconstruction from videos. For example, EndoNeRF [25] follows the modeling of D-NeRF [18] to represent deformable surgical scenes as the combination of a canonical neural radiance field and a time-dependent neural displacement field, and LerPlane [28] factorizes six 2D planes for static field and dynamic field to accelerate optimization similar to [5, 2]. There is also a growing interest to reconstruct surfaces from endoscopic videos via neural implicit fields [31, 1].
@@
Recently 3D Gaussian Splatting (3D-GS) [8] has been witnessed as a powerful representation that renders higher-quality results at a real-time level. Follow-up works [14, 30, 26, 29] extend 3D-GS to represent dynamic scenes and achieve state-of-the-art performances on rendering fidelity and speed. To model dynamic representation of dynamic scenes, [14, 29] formulate 4D Gaussians and assign extra parameters as attributes in the time dimension, and [26, 30] apply MLPs to predict the deformation of the Gaussians, sharing the same rationale with dynamic NeRFs [17, 18]. In this paper, we present EndoGS, a method based on Gaussian Splatting for high-fidelity rendering of deformable endoscopic tissues reconstruction with better rendering quality and better rendering speed.
@@
To summarize, our main contributions are three-fold: 1) We present a novel Gaussian Splatting based method for deformable endoscopic tissues reconstruction. This is one of the first attempts [10] introducing Gaussian Splatting in the medical domain. 2) We represent dynamic surgical scenes with the combination of static Gaussians and the deformable parameters in the time dimension, adopt depth-guided supervision for monocular optimization, and apply a spatial-temporal weight mask to mitigate tool occlusion. Besides, we also introduce total variation items to mitigate quality degradation in the spatial domain as well as in the temporal domain. 3) While former works [25, 28, 27] apply tool masks to filter out un-desired parts and evaluate the rendered visible pixels with the ground-truth, different training masks are adopted for metrics evaluation. To this end, we use the same input masks involved in the training and inference for comparison methods and make a clear and fair comparison on rendered video quality.
@@ 2.1 Overview
In this paper, we introduce our method, referred to as EndoGS, which utilizes a deformable variant of 3D-GS to reconstruct 3D surgical scenes from a single-viewpoint video, estimated depth maps, and labeled tool masks. Specifically, given a stereo video with left and right frames $\{(\bm{I}_{i}^{l},\bm{I}_{i}^{r})\}_{i=1}^{T}$, where $T$ is the total number of frames, our goal is to reconstruct 3D representations of the deformable tissues that render in high quality. We follow the approach of [25, 28, 27] by combining the extracted binary tool masks $\{\bm{M}_{i}\}_{i=1}^{T}$ and the depth maps $\{\bm{D}_{i}\}_{i=1}^{T}$ estimated from binocular captures for the left views. The pipeline of EndoGS is illustrated in Fig. 1. In this section, we first introduce the preliminary of 3D Gaussian Splatting (Section 2.2). We then present the modeling of deformable tissues with a dynamic version of 3D-GS, which adopts a lightweight MLP to represent the dynamic field (Section 2.3). Finally, the training optimization of the Gaussian Splatting with the tool masks and the depth maps is introduced (Section 2.4).
@@ 2.2 Preliminaries of 3D Gaussian Splatting
$G(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\bm{\mu})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x-\bm{\mu}})}.$ (1)
The 3D Gaussians can be projected onto 2D space and rendered for pixels [34]:
$\mathbf{\Sigma}^{\prime}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{T}\mathbf{J}^{T},$ (2)
where $\mathbf{\Sigma}^{\prime}$ is the covariance matrix in the 2D plane, $\mathbf{W}$ denotes view transformation, and the Jacobian $\mathbf{J}$ is the affine approximation of the projective transformation. To enforce the positive semi-definiteness, $\mathbf{\Sigma}$ is parameterized with a scale $\mathbf{S}$ and rotation $\mathbf{R}$:
$\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}.$ (3)
To render the color of the pixels $\mathbf{p}$, point-based volume rendering is adopted:
$C(\mathbf{p})=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (4)
where $\textit{where }\alpha_{i}=\sigma_{i}e^{-\frac{1}{2}(\mathbf{p-\bm{\mu}_{i}})^{T}{\mathbf{\Sigma}^{\prime}}^{-1}(\mathbf{p-\bm{\mu}_{i}})}.$ (5)
The $c_{i}$ means the color of the Gaussians along the ray, $\bm{\mu}_{i}$ denotes the coordinates, and $\sigma_{i}$ denotes the opacity. And 3D-GS applies spherical harmonics [19] to model the view-dependent color. In total, the explicit 3D Gaussians are characterized by: position $\bm{\mu}\in\mathbb{R}^{3}$, scaling factor $\mathbf{s}\in\mathbb{R}^{3}$, rotation factor $\mathbf{r}\in\mathbb{R}^{4}$, spherical harmonic (SH) coefficients $\mathbf{sh}\in\mathbb{R}^{k}$ ($k$ means number of SH functions), and opacity $\sigma\in\mathbb{R}$. Finally each Gaussian can be represented as ($\bm{\mu},\mathbf{s},\mathbf{r},\mathbf{sh},\sigma$).
@@ 2.3 Gaussian Splatting Representations for Deformable Tissues
We represent a surgical scene as a 4D volume, where the deformation of the tissues involves time time dimension. To this end, we use introduce the Gaussian deformation to represent the time-varying motions and shapes, following the basic designs of [26]. The final goal is to learn the original representation of the 3D Gaussians $\{(\bm{\mu},\mathbf{s},\mathbf{r},\mathbf{sh},\sigma)\}$ as well as the Gaussian deformations $\{\Delta(\bm{\mu},\mathbf{s},\mathbf{r},\mathbf{sh},\sigma)\}=\{(\Delta\bm{\mu},\Delta\mathbf{s},\Delta\mathbf{r},\Delta\mathbf{sh},\Delta\sigma)\}$.
@@
$f_{voxel}({\bm{\mu}},t)=\mathcal{B}\left(F_{XY},x,y\right)\odot\mathcal{B}\left(F_{YZ},y,z\right)\ldots\mathcal{B}\left(F_{YT},y,\tau\right)\odot\mathcal{B}\left(F_{ZT},z,\tau\right),$
@@
$\Delta{(\bm{\mu},\mathbf{s},\mathbf{r},\mathbf{sh},\sigma)}={\rm MLP}(f_{voxel}({\bm{\mu}},t)).$
@@ 2.4 Training Combined with Tool Masks and Depth Maps
Reconstructing from videos with tool occlusion poses a challenge and we follow former works [25, 28, 27] to use labeled tool occlusion masks to indicates the unseen pixels. Furthermore, we leverage spatiotemporal importance sampling strategy to indicate the crucial areas related to the occlusion issue. Since we optimize the Gaussians with spatial targets, we can simply devise weight masks in the spatial domain to serve the optimization losses. To be specific, the binary masks are denoted as $\{\bm{M}_{i}\}^{T}_{i=1}$, where 0 stands for tissue pixels and 1 stands for tool pixels, and the importance maps $\{\bm{\mathcal{V}}_{i}\}^{T}_{i=1}$ involve temporal statistics and are denoted as
$\bm{\mathcal{V}}_{i}=(\bm{1}-\bm{M}_{i})\odot\left(\bm{1}+\alpha\sum_{j=1}^{T}\bm{M}_{j}/\left\|\sum_{j=1}^{T}\bm{M}_{j}\right\|_{F}\right).$
We only optimize in the seen part by introducing the item $\bm{1}-\bm{M}_{i}$. Meanwhile, the statistics of the mask frequencies normalized by the Frobenius norm along the temporal dimension provide the information of the uncertainty and allocate higher importance for tissue areas with higher occlusion frequencies. The parameter $\alpha$ is used to control the scaling strength. We use the $L_{1}$ reconstruction loss with spatial masks, and thus the spatial supervision on $i$-th is:
$\mathcal{L}_{L1}(i)=|\bm{I}_{i}\odot\bm{\mathcal{V}}_{i}-\hat{\bm{I}}_{i}\odot\bm{\mathcal{V}}_{i}|,$
@@
Monocular reconstruction provides limited information for 3D reconstruction and makes overfitting happen with single-viewpoint images [4]. To mitigate illness from single-viewpoint inputs, we introduce depth-guided loss with the estimated depth maps. The coarse stereo depth maps $\{\bm{D}_{i}\}^{T}_{i=1}$ are obtained via STTR-light [11]. We adopt Huber loss for depth regularization following [28]:
$\mathcal{L}_{D}(i)=\begin{cases}0.5\Delta\bm{D}_{i}^{2},&\text{ if }|\Delta\bm{D}_{i}|<\delta\\ \delta(\Delta\bm{D}_{i}-0.5\delta),&\text{ otherwise }\end{cases}$
where $\Delta\bm{D}_{i}=|\bm{D}_{i}-\hat{\bm{D}}_{i}|$ denotes the depth loss at $i$-th frame, and $\hat{\bm{D}}_{i}$ is the rendering depth.
@@
$\mathcal{L}(i)=\mathcal{L}_{L1}(i)+\lambda_{D}\mathcal{L}_{D}(i)+\lambda_{TV1}\mathcal{L}_{tv-spatial}(i)+\lambda_{TV2}\mathcal{L}_{tv-temporal},$
where hyperparameters $\lambda_{D}$, $\lambda_{TV1}$, and $\lambda_{TV2}$ control the regularization strength.

### v3

@@ Abstract
Surgical 3D reconstruction is a critical area of research in robotic surgery, with recent works adopting variants of dynamic radiance fields to achieve success in 3D reconstruction of deformable tissues from single-viewpoint videos. However, these methods often suffer from time-consuming optimization or inferior quality, limiting their adoption in downstream tasks. Inspired by 3D Gaussian Splatting, a recent trending 3D representation, we present EndoGS, applying Gaussian Splatting for deformable endoscopic tissue reconstruction. Specifically, our approach incorporates deformation fields to handle dynamic scenes, depth-guided supervision with spatial-temporal weight masks to optimize 3D targets with tool occlusion from a single viewpoint, and surface-aligned regularization terms to capture the much better geometry. As a result, EndoGS reconstructs and renders high-quality deformable endoscopic tissues from a single-viewpoint video, estimated depth maps, and labeled tool masks. Experiments on DaVinci robotic surgery videos demonstrate that EndoGS achieves superior rendering quality. Code is available at https://github.com/HKU-MedAI/EndoGS.
@@ 1 Introduction
Reconstruction of high-quality deformable tissues from endoscopic videos is a significant but challenging task, facilitating downstream tasks like surgical AR, education, and robot learning [4, 26, 27]. Earlier attempts [18, 24, 28, 39, 40] adopt depth estimation to achieve great success in endoscopic reconstruction, but these methods still struggle to produce realistic 3D reconstruction due to two key issues. First, non-rigid deformations with sometimes large movements, which require practical dynamic scene reconstruction, hinders the adaptation of those techniques. Second, tool occlusion happens in single-viewpoint videos, producing difficulties in learning affected parts with limited information.
@@
While [12, 17] have proposed frameworks combining tool masking, stereo depth estimation, and sparse warp field [7, 21] for single-viewpoint 3D reconstruction, they are still prone to failure in the presence of dramatic non-topological deformable tissues changes. Neural Radiance Fields (NeRFs) [20] have shown great potentials in 3D representations, and methods based on variants of dynamic radiance field [23, 6] have become representative works in deformable tissues reconstruction from videos. For example, EndoNeRF [29] follows the modeling of D-NeRF [23] to represent deformable surgical scenes as the combination of a canonical neural radiance field and a time-dependent neural displacement field, and LerPlane [33] factorizes six 2D planes for static field and dynamic field to accelerate optimization similar to [6, 2].
@@
Recently 3D Gaussian Splatting (3D-GS) [10, 31, 3] has been witnessed as a powerful representation that renders higher-quality results at a real-time level. Follow-up works [19, 37, 30, 35] extend 3D-GS to represent dynamic scenes and achieve state-of-the-art performances on rendering fidelity and speed. To model dynamic representation of dynamic scenes, [19, 35] formulate 4D Gaussians and assign extra parameters as attributes in the time dimension, and [30, 37] apply MLPs to predict the deformation of the Gaussians, sharing the same rationale with dynamic NeRFs [22, 23]. In this paper, we present EndoGS, a method based on surface-aligned Gaussian Splatting for deformable endoscopic tissues reconstruction with better rendering quality and better rendering speed.
@@
To summarize, our main contributions are three-fold: 1) We present the first Gaussian Splatting based method for deformable endoscopic tissues reconstruction. This is one of the first attempts [13] introducing Gaussian Splatting in the medical domain. 2) We represent dynamic surgical scenes with the combination of static Gaussians and the deformable parameters in the time dimension, adopt depth-guided supervision with spatial-temporal weight masks for monocular optimization with tool occlusion, and combine surface-aligned regularization terms  [8] to capture the much better geometry. 3) We use the same input tool masks involved in the training and inference for comparison methods and make a clear and fair comparison, and experiments demonstrate our superior performances.
@@ 2.1 Overview
In this paper, we introduce our method, referred to as EndoGS, which utilizes a deformable variant of 3D-GS to reconstruct 3D surgical scenes from a single-viewpoint video, estimated depth maps, and labeled tool masks. Specifically, given a stereo video with left and right frames $\{(\bm{I}_{i}^{l},\bm{I}_{i}^{r})\}_{i=1}^{T}$, where $T$ is the total number of frames, our goal is to reconstruct 3D representations of the deformable tissues that render in high quality. We follow the approach of [29, 33, 32] by combining the extracted binary tool masks $\{\bm{M}_{i}\}_{i=1}^{T}$ and the depth maps $\{\bm{D}_{i}\}_{i=1}^{T}$ estimated from binocular captures for the left views. The pipeline of EndoGS is illustrated in Fig. 1. In this section, we first introduce the preliminary of 3D Gaussian Splatting (Section 2.2). We then present the modeling of deformable tissues with a dynamic version of 3D-GS, which adopts a lightweight MLP to represent the dynamic field (Section 2.3). The training optimization of the Gaussian Splatting with the tool masks and the depth maps is introduced (Section 2.4). Finally, we discuss how to align Gaussians with the Surface (Section 2.5).
@@ 2.2 Preliminaries of 3D Gaussian Splatting
$G(\bm{x})=e^{-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\mathbf{\Sigma}^{-1}(\bm{x-\bm{\mu}})}.$ (1)
The 3D Gaussians can be projected onto 2D space and rendered for pixels: $\mathbf{\Sigma}^{\prime}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{T}\mathbf{J}^{T}$, where $\mathbf{\Sigma}^{\prime}$ is the covariance matrix in the 2D plane, $\mathbf{W}$ denotes view transformation, and the Jacobian $\mathbf{J}$ is the affine approximation of the projective transformation. To enforce the positive semi-definiteness, $\mathbf{\Sigma}$ is parameterized as $\mathbf{R}$: $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$ with a scale $\mathbf{S}$ and rotation.
To render the color of the pixels $\bm{p}$, point-based volume rendering is adopted:
$C(\bm{p})=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (2)
where $\textit{where }\alpha_{i}=\sigma_{i}e^{-\frac{1}{2}(\bm{p-\bm{\mu}_{i}})^{T}{\mathbf{\Sigma}^{\prime}}^{-1}(\bm{p-\bm{\mu}_{i}})}.$ (3)
The $c_{i}$ means the color of the Gaussians along the ray, $\bm{\mu}_{i}$ denotes the coordinates, and $\sigma_{i}$ denotes the opacity. And 3D-GS applies spherical harmonics to model the view-dependent color.
In total, the explicit 3D Gaussians are characterized by: position $\bm{\mu}\in\mathbb{R}^{3}$, scaling factor $\bm{s}\in\mathbb{R}^{3}$, rotation factor $\bm{r}\in\mathbb{R}^{4}$, spherical harmonic (SH) coefficients $\bm{sh}\in\mathbb{R}^{k}$ ($k$ means number of SH functions), and opacity $\sigma\in\mathbb{R}$. Finally each Gaussian can be represented as ($\bm{\mu},\bm{s},\bm{r},\bm{sh},\sigma$).
@@ 2.3 Gaussian Splatting Representations for Deformable Tissues
We represent a surgical scene as a 4D volume, where the deformation of the tissues involves time dimension. To this end, we introduce the Gaussian deformation to represent the time-varying motions and shapes, following the basic designs of [30]. The final goal is to learn the original representation of the 3D Gaussians $\{(\bm{\mu},\bm{s},\bm{r},\bm{sh},\sigma)\}$ as well as the Gaussian deformations $\{\Delta(\bm{\mu},\bm{s},\bm{r},\bm{sh},\sigma)\}=\{(\Delta\bm{\mu},\Delta\bm{s},\Delta\mathbf{r},\Delta\bm{sh},\Delta\sigma)\}$.
@@
$f_{voxel}({\bm{\mu}},t)=\mathcal{B}\left(F_{XY},x,y\right)\odot\mathcal{B}\left(F_{YZ},y,z\right)\ldots\mathcal{B}\left(F_{YT},y,t\right)\odot\mathcal{B}\left(F_{ZT},z,t\right).$ (4)
@@
$\Delta{(\bm{\mu},\bm{s},\bm{r},\bm{sh},\sigma)}={\rm MLP}(f_{voxel}({\bm{\mu}},t)).$ (5)
@@ 2.4 Training Combined with Tool Masks and Depth Maps
Reconstructing from videos with tool occlusion poses a challenge and we follow former works [29, 33, 32] to use labeled tool occlusion masks to indicates the unseen pixels. Furthermore, we leverage spatiotemporal importance sampling strategy to indicate the crucial areas related to the occlusion issue. To be specific, the binary masks are denoted as $\{\bm{M}_{i}\}^{T}_{i=1}$, where 0 stands for tissue pixels and 1 stands for tool pixels, and the importance maps $\{\bm{\mathcal{V}}_{i}\}^{T}_{i=1}$ involve temporal statistics and are denoted as $\bm{\mathcal{V}}_{i}=(\bm{1}-\bm{M}_{i})\odot\left(\bm{1}+\alpha\sum_{j=1}^{T}\bm{M}_{j}/\left\|\sum_{j=1}^{T}\bm{M}_{j}\right\|_{F}\right).$ We only optimize in the seen part by introducing the item $\bm{1}-\bm{M}_{i}$. Meanwhile, the statistics of the mask frequencies normalized by the Frobenius norm along the temporal dimension provide the information of the uncertainty and allocate higher importance for tissue areas with higher occlusion frequencies. The parameter $\alpha$ is used to control the scaling strength. We use the $L_{1}$ loss with spatial masks, and thus the spatial supervision on $i$-th is
$\mathcal{L}_{L1}(i)=|\bm{I}_{i}\odot\bm{\mathcal{V}}_{i}-\hat{\bm{I}}_{i}\odot\bm{\mathcal{V}}_{i}|,$ (6)
@@
Monocular reconstruction provides limited information for 3D reconstruction and makes overfitting happen with single-viewpoint images [5]. To mitigate illness from single-viewpoint inputs, we introduce depth-guided loss with the estimated depth maps. The coarse stereo depth maps are obtained via STTR-light [14]. We adopt Huber loss $\mathcal{L}_{D}(i)$ for depth regularization following [33].
@@
2.5 Surface-Aligned Gaussians
When reconstructing the region with limited 3D cues, especially for surrounding tool occlusion masks, noticeable artifacts on the surface are evident. [8] proposes a technique aimed at facilitating precise and rapid mesh extraction from 3D-GS.
To ensure tight integration of the Gaussians with the surface, surface alignment normalization is applied by controlling the density function of the Gaussians, which can be defined as $d(\bm{p})=\sum_{g}\sigma_{g}\exp\left(-\frac{1}{2}(\bm{p}-\bm{\mu}_{g})^{T}\mathbf{\Sigma}^{-1}_{g}(\bm{p}-\bm{\mu}_{g})\right)\>$, where $\bm{p}$ denotes the position and $g$ denotes the Gaussian.
If the Gaussians are aligned with the surface, we could have three assumptions. 1) The closest Gaussian $g^{*}=\arg\min_{g}\left\{(\bm{p}-\bm{\mu}_{g})^{T}\mathbf{\Sigma}^{-1}_{g}(\bm{p}-\bm{\mu}_{g})\right\}\>$ to the point $\bm{p}$ is likely to contribute much more than others to the density value $d(\bm{p})$. 2) To ensure the 3D Gaussians to be flat, every Gaussian $g$ would have one of its three scaling factors close to zero. 3) Gaussians are opaque and we can encourage $\sigma_{i}$ to be 1 with cross-entropy loss. Under these assumptions, the density can be approximated as $\bar{d}(\bm{p})=\exp\left(-\frac{1}{2s_{g^{*}}^{2}}\langle\bm{p}-\bm{\mu}_{g^{*}},\bm{n}_{g^{*}}\rangle^{2}\right)\>$, where $s_{g^{*}}$ is the smallest value of $\bm{s}_{g^{*}}$ and $\bm{n}_{g^{*}}$ is the direction of the corresponding axis. Based on the approximated expression of $\bar{d}(\bm{p})$, we could find the zero-crossings of the Signed Distance Function and approximate the ideal distance function associated with the density function $d$ as $f(\bm{p})=\pm s_{g*}\sqrt{-2\log\left(d(\bm{p})\right)}.$ We do not use the density loss in the implementation of [8] and refer to [8] for details of the approximation.
We use the regularization term on the ideal SDF $f(\bm{p})$ and an estimated $\hat{f}(\bm{p})$ that is the difference between the depth of $\bm{p}$ and the depth in the corresponding depth map at the projection of $\bm{p}$:
$\mathcal{L}_{SDF}=\frac{1}{|\mathcal{P}|}\sum_{\bm{p}\in\mathcal{P}}|\hat{f}(\bm{p})-f(\bm{p})|.$ (7)
The estimated $\hat{f}(\bm{p})$ is defined as the difference between the depth of $\bm{p}$ and the depth in the corresponding depth map at the projection of $\bm{p}$. Meanwhile, we add a regularization term for normal direction:
$\mathcal{L}_{norm}=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}\left\|\frac{\nabla f(\bm{p})}{\|\nabla f(\bm{p})\|_{2}}-\bm{n}_{g^{*}}\right\|_{2}^{2}\>.$ (8)
Besides, to encourage opaque Gaussians, we adopt cross-entropy loss for opacity regularization:
$\mathcal{L}_{opacity}(i)=-\sum_{j}(\sigma+\Delta\sigma)_{j}\log(\sigma+\Delta\sigma)_{j},$ (9)
where $(\sigma+\Delta\sigma)_{j}$ denotes the $j$-th Gaussian’s opacity in the $i$-th frame.
To sum up, our final optimization target at the $i$-th frame is:
$\mathcal{L}(i)=\mathcal{L}_{L1}(i)+\lambda_{D}\mathcal{L}_{D}(i)+\lambda_{TV1}\mathcal{L}_{tv-spatial}(i)+\lambda_{TV2}\mathcal{L}_{tv-temporal}+\lambda_{S}\mathcal{L}_{S}(i),$ (10)
where $\mathcal{L}_{S}(i)=\mathcal{L}_{SDF}+\mathcal{L}_{norm}+0.5\mathcal{L}_{opacity}(i)$ denotes the surface-aligned regularization. Hyperparameters $\lambda_{D}$, $\lambda_{TV1}$, $\lambda_{TV2}$, $\lambda_{S}$ control the strength.